OpenStudy (anonymous):

Solve each system using substitution. Check your answer. 3x+2y=23 1/2x-4=y

Mehek (mehek14):

didn't I just solve this?

OpenStudy (anonymous):

First, plug 1/2x-4 into 3x+2y=23. Solve for x. Then plug the x value you get into 3x+2y=23. Your answers should be x=7.75, y=-1/8.